OpenStudy (anonymous):

Solve the equation: x2 − 9x + 23 = 0

OpenStudy (anonymous):

one way to do this is to use the formula x = [-b +-sqrt(b^2-4ac)] / 2a where the equation is a x^2 + bx + c = 0 so for your equation you plug in a=1 , b = -9 and c = 23 into the above formula

OpenStudy (anonymous):

\[\left\{x=\frac{1}{2} \left(9-i \sqrt{11}\right),x=\frac{1}{2} \left(9+i \sqrt{11}\right)\right\} \]